OpenStudy (anonymous):

A local store owner pays her employees time and a half for overtime. That means for every hour an employee works more than 40 hours per week, the store will pay 1.5 times the regular hourly wage of $8.50. Write a function o(x) that defines the hourly pay for overtime hours worked.

OpenStudy (anonymous):

i confused

OpenStudy (anonymous):

was the other one wrong?

OpenStudy (texaschic101):

I am sorry, I am not good at this.....time and a half for 8.75 =13.13 f(x) = 13.13(x) time and a half for 8.75 = 13.13 per overtime hour

OpenStudy (texaschic101):

you might better get a second opinion.....

OpenStudy (anonymous):

at least your trying to help, but how did u get 13.13?

OpenStudy (texaschic101):

time and a half for 8.75 = 1.5 x 8.75 = 13.125 rounded to 13.13

OpenStudy (texaschic101):

oh wait...it is 8.50....let me change it

OpenStudy (anonymous):

i think u where right the first time with 12.75

OpenStudy (texaschic101):

f(x) = 12.75(x) 12.75 is time and a half for 8.50 I guess I was ...lol

OpenStudy (anonymous):

are you sure about this one?

OpenStudy (texaschic101):

I think so.....you still might need a second opinion....what do you think ?

OpenStudy (anonymous):

do u know how to write a piecewise function? I'll ask just in case.

OpenStudy (texaschic101):

lol.....no sorry....not good at functions

OpenStudy (anonymous):

me either.

jimthompson5910 (jim_thompson5910):

we're told to "Write a function o(x) that defines the hourly pay for overtime hours worked." so because overtime only kicks in when you work "more than 40 hours per week", this means that x > 40

jimthompson5910 (jim_thompson5910):

the overtime pay is "1.5 times the regular hourly wage of $8.50" so the overtime hourly pay is 1.5*8.5 = 12.75 dollars an hour

OpenStudy (anonymous):

so u multiply 1.5 and 8.5.

jimthompson5910 (jim_thompson5910):

so if x is some number over 40, then x - 40 is the number of hours you work overtime (eg: if x = 50, then x-40 = 50 - 40 = 10 and you worked 10 overtime hours) so if you want to figure out how much overtime you earn, you just multiply the two to get 12.75(x-40) optionally you can distribute, but it's fine to leave it like that

OpenStudy (anonymous):

got it. do u know how to write a piecewise function?

jimthompson5910 (jim_thompson5910):

yes I do

OpenStudy (anonymous):

can u help me with this question? combine the two function you just wrote to form a piewise function, W(x), in which each function is defined for a specific domain. so H(x)=8.5x and O(x)=12.75(x-40)

jimthompson5910 (jim_thompson5910):

a piecewise function is just a combination of two or more functions

jimthompson5910 (jim_thompson5910):

what you do is basically write all the functions in one column next to each function, you describe the conditions or restrictions on each piece of that piecewise function

OpenStudy (anonymous):

i dont now how to organize the function.

jimthompson5910 (jim_thompson5910):

so say f(x) is the piecewise function

jimthompson5910 (jim_thompson5910):

we would then say f(x) = 8.5x if x<= 40 OR f(x) = 12.75(x-40) if x>40

jimthompson5910 (jim_thompson5910):

notice how x dictates which function to use if x <= 40, then you use the first piece if x > 40, then you go for the second piece you can only use one piece at a time

OpenStudy (anonymous):

so it doesnt have to do with greater or less.

jimthompson5910 (jim_thompson5910):

you would then combine the two and write it in this notation